# thanks: The first two authors contributed equally

### v1

@@ II.4 Parameter estimation setup
We obtain our results by performing Bayesian parameter inference runs using state-of-the-art GW parameter inference framework built in Bilby Ashton et al. (2019) with a dynamical nested sampler dynesty Speagle (2020). For the injection and recovery, we use the publicly available eccentric waveform model TEOBResumS-Dalí Nagar et al. (2018, 2021)  [^5] We use the commit 0f19532 of the eccentric branch of TEOBResumS available at https://bitbucket.org/eob_ihes/teobresums/. In the following, we will denote the eccentricity that the waveform model takes as an input by $e_{\text{eob}}$ and the frequency where $e_{\text{eob}}$ is defined by $f_{\text{ref}}$. Additionally, we provide $\langle f_{\text{ref}}\rangle,e_{\text{gw}}$} for each injection to facilitate better comparisons across waveform models and systematics studies.  [^6] Note that we sample over the model eccentricity $e_{\text{eob}}$ and not $e_{\text{gw}}$.
Ideally, in order to accurately describe an eccentric orbit, the waveform model should also include mean anomaly as a free parameter Islam et al. (2021); Clarke et al. (2022); Shaikh et al. (2023). Most of the publicly available eccentric waveform models like TEOBResumS-Dalí, SEOBNRE Cao and Han (2017); Liu et al. (2020), EccentricTD Tanay et al. (2016) do not include mean anomaly as a free parameter but instead start at the pericenter or apocenter.  [^7] Ecc1dSur Islam et al. (2021) and SEOBNRv4EHM Ramos-Buades et al. (2022b, 2023) include mean anomaly as a free parameter but not yet publicly available. This, however, does not affect our results since we use the same waveform model to inject and recover.
@@
To obtain the 2D posteriors on $\Delta M_{f}/\bar{M}_{f}$ and $\Delta\chi_{f}/\bar{\chi}_{f}$ for the IMRCT, we use the summarytgr command line tool from PESummary Hoy and Raymond (2021). The uninformative priors used for the mass and spin parameters lead to non-trivial priors on $\Delta M_{f}/\bar{M}_{f}$ and $\Delta\chi_{f}/\bar{\chi}_{f}$. To impose agnostic priors on these inference parameters used to probe GR-violation, we reweight the 2D posterior to a flat prior on the deviation parameters, as is routinely done by the LVK Abbott et al. (2021c, b).
@@
Given the posterior samples obtained in the parameter estimation from the inspiral and post-inspiral part, summarytgr uses an NR informed fit to obtain the final mass and spin posteriors. This fit is based on quasicircular NR simulations only and does not take into account any eccentric corrections which has been shown to be negligible at the quasi-Newtonian level with small eccentricity approximation Bhat et al. (2023). Currently, we lack a sufficient number of NR-simulations of merging eccentric binaries required for obtaining the eccentricity corrected NR-fit for final mass and final spin. We measure $\langle f_{\text{ref}}\rangle,e_{\text{gw}}$} for a given {$f_{\text{ref}},e_{\text{eob}}$} using gw_eccentricity Shaikh et al. (2023).
@@ III Results
The main result of this work can be divided into two parts. The first part, Sec. III.1, discusses the effect of neglecting eccentricity in the waveform template on the IMRCT. The second part, Sec. III.2, demonstrates the robustness of the IMRCT performed on GW data containing an eccentric signal using a waveform template that includes physics required to describe GWs coming from binaries on an eccentric orbit. We summarize the results of the study of bias in IMRCT due to neglecting eccentricity in Table 2.
@@ III.1.1 Aligned spin case
Since TEOBResumS-Dalí can describe only non-precessing systems, $\theta^{\mathbf{S}\mathbf{L}}_{i}$ can take only two values 00 (aligned) or $\pi$ (antialigned). In this section, we consider binary systems with $\theta^{\mathbf{S}\mathbf{L}}_{1,2}=0$.
We choose the component masses to be similar to that of the gravitational wave event GW150914. For this study we choose a total mass of system that lies within the ‘golden mass’ range $M\in[50-100]M_{\odot}$ which ensures high enough SNRs both in the inspiral and post-inspiral parts of the signal Ghosh et al. (2016). The injection parameters and the priors are noted in Table 1. The injection eccentricities $e_{\text{eob}}\in$0.05, 0.06, 0.07, 0.08, 0.09, 0.1} are defined at a reference frequency of $f_{\text{ref}}=$ 20 Hz. These injected waveforms have $e_{\text{gw}}$ values {0.040, 0.048, 0.055, 0.063, 0.071, 0.079}  [^8] We used the ResidualAmplitude method within gw_eccentricity to measure these eccentricities. measured at an orbit averaged frequency of $\langle f_{\text{ref}}\rangle=$ 25 Hz. To understand the effect of spin magnitudes on IMRCT systematics, we consider two different choices of dimensionless component spins, one low $(\chi_{1},\chi_{2})=(0.2,0.1)$ and one moderate $(\chi_{1},\chi_{2})=(0.4,0.3)$, for the set of injected eccentricities. All the injections considered in this study have HLV network (at O4 design sensitivity) signal to noise ratio (SNR) in the range $27-52$.
@@
Table 1: Injection values of the binary parameters and corresponding priors used for parameter estimation with Bilby Ashton et al. (2019). $\mathcal{U}(a,b)$ stands for uniform prior in the range $(a,b)$. $\mathcal{\sin}(0,\pi)$ and $\mathcal{\cos}(-\pi/2,\pi/2)$ stand for the Sine and Cosine priors, respectively on the ranges mentioned. Single values represent Dirac delta priors peaked at those values. Different values in the the curly braces represent different injections. The table, therefore, represents 12 injections – 6 different eccentricities each with 2 different spin configurations.
The results of these IMRCTs are shown in Fig. 1. The top row shows the results for $(\chi_{1},\chi_{2})=(0.4,0.3)$ and the bottom row shows the same for $(\chi_{1},\chi_{2})=(0.2,0.1)$. In each of these rows, the left panel shows the iso-probability contours of the 2D probability distribution on the final mass deviation $\Delta M_{f}/\bar{M}_{f}$ (along $x$-axis) and the final spin deviation $\Delta\chi_{f}/\bar{\chi}_{f}$ (along $y$-axis) plane at 90% confidence. The right panel shows the same at 68% confidence. Different contours in each panel correspond to the injected eccentricities $e_{\text{eob}}\in$0.05, 0.06, 0.07, 0.08, 0.09, 0.1} as noted on the top of the right panel. Fig. 1 also shows the marginalised 1D posterior of $\Delta M_{f}/\bar{M}_{f}$ on top of the left panel and that of $\Delta\chi_{f}/\bar{\chi}_{f}$ on the right of the right panel.
@@ III.1.2 Antialigned spin case
    $(\pi,0)$$\mathbf{S}_{1}$ is antialigned and $\mathbf{S}_{2}$ is aligned.
@@
    $(0,\pi)$$\mathbf{S}_{1}$ is aligned and $\mathbf{S}_{2}$ is antialigned.
@@
    $(\pi,\pi)$ – Both $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ are antialigned.
@@ Acknowledgements.
We thank Pankaj Saini for a careful reading of the draft. We also thank Prayush Kumar, Nathan K. Johnson-McDaniel, Aditya Vijaykumar, and members of the ICTS-TIFR Astrophysical Relativity Group, for useful discussions. M.A.S.’s research was supported by the Department of Atomic Energy, Government of India under project no. RTI4001 and the National Research Foundation of Korea under grant No. NRF-2021R1A2C2012473. S.A.B. acknowledges support from the Department of Science and Technology and the Science and Engineering Research Board (SERB) of India via Swarnajayanti Fellowship Grant No. DST/SJF/PSA-01/2017-18 and support from Infosys foundation. The authors acknowledge the public code at https://github.com/osheamonn/eobresum_bilby/, a modified version of which was used as the source model for injection study with Bilby Ashton et al. (2019). The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. The authors also acknowledge the use of the Alice cluster at the International Centre for Theoretical Sciences.

### v2

@@ II.4 Parameter Estimation setup
We obtain our results by performing Bayesian parameter inference runs using state-of-the-art GW parameter inference framework built in Bilby Ashton et al. (2019) with a dynamical nested sampler dynesty Speagle (2020). For the injection and recovery, we use the publicly available eccentric waveform model TEOBResumS-Dalí Nagar et al. (2018, 2021b); Chiaramello and Nagar (2020); Albanesi et al. (2021); Nagar et al. (2021a); Nagar and Rettegno (2021); Nagar et al. (2024)  [^6] We use the commit 0f19532 of the eccentric branch of TEOBResumS available at https://bitbucket.org/eob_ihes/teobresums/. In the following, we will denote the eccentricity that the waveform model takes as an input by $e_{\text{eob}}$ and the frequency where $e_{\text{eob}}$ is defined by $f_{\text{ref}}$. Additionally, we provide $\langle f_{\text{ref}}\rangle,e_{\text{gw}}$} for each injection to facilitate better comparisons across waveform models and systematics studies.  [^7] Note that we sample over the model eccentricity $e_{\text{eob}}$ and not $e_{\text{gw}}$. For a given model eccentricity $e_{\text{eob}}$ at a model reference frequency $f_{\text{ref}}$, the standardized eccentricity $e_{\text{gw}}$ at $\langle f_{\text{ref}}\rangle$ is obtained by first generating the eccentric waveform modes for the given pair $\{f_{\text{ref}},e_{\text{eob}}\}$ and then following the steps discussed in Sec. II.3.
Ideally, to accurately describe an eccentric orbit, the waveform model should also include mean anomaly as a free parameter Islam et al. (2021); Clarke et al. (2022); Shaikh et al. (2023). In this work, we use the version of TEOBResumS-Dalí that was used in Ref. Bonino et al. (2023) by the waveform developers to showcase the accuracy and robustness of their model. This version of TEOBResumS-Dalí does not allow mean anomaly as a free parameter. Although the most recent version of TEOBResumS-Dalí, released after the completion of this work, allows mean anomaly as a free parameter, this version is yet to be tested (at the time of writing) in a parameter inference study by sampling over mean anomaly and is currently under review. Eccentric waveform models Ecc1dSur Islam et al. (2021) and SEOBNRv4EHM Ramos-Buades et al. (2022a, 2023) include mean anomaly as a free parameter but these are not publicly available. Publicly available eccentric waveform models SEOBNRE Cao and Han (2017); Liu et al. (2020) and EccentricTD Tanay et al. (2016) do not include mean anomaly as a free parameter but instead start at the pericenter or apocenter. [^8] In some classes of GW sources, eccentricity and mean anomaly may not be sufficient to describe the orbit. For example, in Ref. Carullo et al. (2024), the Arnowitt-Deser-Misner energy and angular momentum are used to describe a generic planar orbit. See also Ref. Gamba et al. (2023) which use the same parameterization to study the possibility of GW190521 resulting from a dynamical capture of two nonspinning black holes.
Because the version of TEOBResumS-Dalí used in the work does not allow sampling over mean anomaly, Ref. Bonino et al. (2023) sampled over the reference frequency, i.e., the PE was performed by sampling over eccentricity and reference frequency with fixed mean anomaly. The motivation behind such an approach is that by varying the reference frequency with fixed mean anomaly one can expect to explore the same parameter space as in the case of varying mean anomaly at a fixed reference frequency. For example, in a recent work in Ref. Bonino et al. (2024), it was shown that one can map an eccentric Numerical Relativity (NR) waveform to an eccentric effective one body waveform using just the initial eccentricity and reference frequency at the first apocenter (that is, at a fixed mean anomaly). Although sampling over reference frequency may reduce bias in PE compared to a fixed reference frequency when mean anomaly is fixed, no significant bias was found in the recovered parameters when sampling exclusively over eccentricity with both mean anomaly and reference frequency fixed (see Fig. 3. in Ref. Bonino et al. (2023)).
Contrary to Ref. Bonino et al. (2023, 2024) which suggest that using (eccentricity, mean anomaly) at fixed reference frequency and using (eccentricity, reference frequency) at fixed mean anomaly is equivalent, it has also been argued that sampling over eccentricity and reference frequency with fixed mean anomaly may not be equivalent to sampling over eccentricity and mean anomaly with a fixed reference frequency (see the introduction in Ref. Gupte et al. (2024)) since these two methods may not explore the same parameter space. In addition, varying reference frequency changes the reference point where binary properties are measured.
Given the fact that no significant bias was reported by the waveform developers in Ref. Bonino et al. (2023) when sampling only eccentricity with fixed reference frequency and that this version does not allow varying mean anomaly, we restrict our PE to sample only over eccentricity. However, future works that will use the recent version of TEOBResumS-Dalí once it is reviewed, should sample over mean anomaly to reduce any potential bias due to neglecting the mean anomaly.
@@
To obtain the 2D posteriors on $\Delta M_{f}/\bar{M}_{f}$ and $\Delta\chi_{f}/\bar{\chi}_{f}$ for the IMRCT, we use the summarytgr command line tool from PESummary Hoy and Raymond (2021). It provides the posteriors on $\Delta M_{f}/\bar{M}_{f}$ and $\Delta\chi_{f}/\bar{\chi}_{f}$ in a post-processing step making use of mass and spin posterior samples along with the NR-fits for the final mass and the final spin parameters. The uninformative priors used for the mass and spin parameters force non-trivial undesirable priors on $\Delta M_{f}/\bar{M}_{f}$ and $\Delta\chi_{f}/\bar{\chi}_{f}$ obtained in the post-processing. To revoke this effect and to impose agnostic priors on these inference parameters used to probe GR-violation, we reweight [^9] Reweighting of posterior refers to removing the effect of any non-trivial undesirable prior on the posterior (see Chapter 5 in Ref. Gelman (2013) for details). This is done when there is no prior information about the parameter in consideration or the prior information about the parameter is unphysical and one wants to be as objective as possible. the 2D posterior to a flat prior on the deviation parameters, as is routinely done by the LVK Abbott et al. (2021c, b).
@@
Given the posterior samples obtained in the PE from the inspiral and post-inspiral part, summarytgr uses an NR informed fit to obtain the final mass and spin posteriors. This fit is based on quasicircular NR simulations only and does not take into account any eccentric corrections which has been shown to be negligible at the quasi-Newtonian level with small eccentricity approximation Bhat et al. (2023). Currently, we lack a sufficient number of NR-simulations of merging eccentric binaries required for obtaining the eccentricity corrected NR-fit for final mass and final spin. We measure $\langle f_{\text{ref}}\rangle,e_{\text{gw}}$} for a given {$f_{\text{ref}},e_{\text{eob}}$} using gw_eccentricity Shaikh et al. (2023).
@@ III Results
The main result of this work can be divided into two parts. The first part, Sec. III.1, discusses the effect of neglecting eccentricity in the waveform template on the IMRCT. The second part, Sec. III.2, demonstrates the robustness of the IMRCT performed on GW data containing an eccentric signal using a waveform template that includes physics required to describe GWs coming from binaries on an eccentric orbit. We list the injection values and priors in 1 and summarize the results of the study of bias in IMRCT due to neglecting eccentricity in 2.
@@ III.1.1 Aligned spin case
Since TEOBResumS-Dalí can describe only non-precessing systems, $\theta^{\mathbf{S}\mathbf{L}}_{i}$ can take only two values 00 (aligned) or $\pi$ (antialigned). In this section, we consider binary systems with $\theta^{\mathbf{S}\mathbf{L}}_{1,2}=0$. Since the waveform model only allows tilts of either 0 or $\pi$, and does not allow sampling over arbitrary spin orientations, we choose the prior on the tilts to be a Dirac delta function peaked at the injection value.
We choose the component masses to be similar to that of the gravitational wave event GW150914. For this study we choose a total mass of system that lies within the ‘golden mass’ range $M\in[50-100]M_{\odot}$ which ensures high enough SNRs both in the inspiral and post-inspiral parts of the signal Ghosh et al. (2016). The injection parameters and the priors are noted in Table 1. The injection eccentricities $e_{\text{eob}}\in$0.05, 0.06, 0.07, 0.08, 0.09, 0.1} are defined at a reference frequency of $f_{\text{ref}}=$ 20 Hz. These injected waveforms have $e_{\text{gw}}$ values {0.040, 0.048, 0.055, 0.063, 0.071, 0.079}  [^10] We used the ResidualAmplitude method within gw_eccentricity to measure these eccentricities. measured at an orbit averaged frequency of $\langle f_{\text{ref}}\rangle=$ 25 Hz. To understand the effect of spin magnitudes on IMRCT systematics, we consider two different choices of dimensionless component spins, one low $(\chi_{1},\chi_{2})=(0.2,0.1)$ and one moderate $(\chi_{1},\chi_{2})=(0.4,0.3)$, for the set of injected eccentricities. All the injections considered in this study have Hanford-Livingston-Virgo network (at O4 design sensitivity) signal to noise ratio (SNR) in the range $27-52$.
@@
Table 1: Injection values of the binary parameters and corresponding priors used for PE with Bilby Ashton et al. (2019). $\mathcal{U}(a,b)$ stands for uniform prior in the range $(a,b)$. $\mathcal{\sin}(0,\pi)$ and $\mathcal{\cos}(-\pi/2,\pi/2)$ stand for the Sine and Cosine priors, respectively on the ranges mentioned. Single values represent Dirac delta priors peaked at those values. Different values in the the curly braces represent different injections. The table, therefore, represents 12 injections – 6 different eccentricities each with 2 different spin configurations.
The results of these IMRCTs are shown in Fig. 1. The top row shows the results for $(\chi_{1},\chi_{2})=(0.4,0.3)$ and the bottom row shows the same for $(\chi_{1},\chi_{2})=(0.2,0.1)$. In each of these rows, the left panel shows the iso-probability contours of the 2D probability distribution on the final mass deviation $\Delta M_{f}/\bar{M}_{f}$ (along $x$-axis) and the final spin deviation $\Delta\chi_{f}/\bar{\chi}_{f}$ (along $y$-axis) plane at 90% confidence. The right panel shows the same at 68% confidence. Different contours in each panel correspond to the injected eccentricities $e_{\text{eob}}\in$0.05, 0.06, 0.07, 0.08, 0.09, 0.1} as noted on the top of the right panel. Fig. 1 also shows the marginalized 1D posterior of $\Delta M_{f}/\bar{M}_{f}$ on top of the left panel and that of $\Delta\chi_{f}/\bar{\chi}_{f}$ on the right of the right panel.
@@ III.1.2 Antialigned spin case
    $(\pi,0)$: $\mathbf{S}_{1}$ is antialigned and $\mathbf{S}_{2}$ is aligned.
@@
    $(0,\pi)$: $\mathbf{S}_{1}$ is aligned and $\mathbf{S}_{2}$ is antialigned.
@@
    $(\pi,\pi)$: Both $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ are antialigned.
@@ Acknowledgements.
We thank the anonymous referees for their useful comments and suggestions. We thank Pankaj Saini for a careful reading of the draft. We also thank Prayush Kumar, Nathan K. Johnson-McDaniel, Aditya Vijaykumar, and members of the ICTS-TIFR Astrophysical Relativity Group, for useful discussions. M.A.S.’s research was supported by the Department of Atomic Energy, Government of India under project no. RTI4001 and the National Research Foundation of Korea under grant No. NRF-2021R1A2C2012473. S.A.B. acknowledges support from the Department of Science and Technology and the Science and Engineering Research Board (SERB) of India via Swarnajayanti Fellowship Grant No. DST/SJF/PSA-01/2017-18 and support from Infosys foundation. The authors acknowledge the public code at https://github.com/osheamonn/eobresum_bilby/, a modified version of which was used as the source model for injection study with Bilby Ashton et al. (2019). The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. The authors also acknowledge the use of the Alice cluster at the International Centre for Theoretical Sciences.